# Extending Linear Convergence of the Proximal Point Algorithm: The Quasar-Convex Case

### v1

@@ 5 Numerical Experiments
In the examples below, we illustrate our results for classes of functions which are neither convex nor quasiconvex, since their sublevel sets are not convex, which represent a hard problem for first-order algorithms.
@@ 5.1 Implementations
To solve the above two examples in this section via PPA, we can not use the common smooth first-order method or second-order method for the subsproblems since both subproblems involved in Examples 27 and 28 are nonsmooth. Therefore, we use the Clarke necessary optimality condition [11, Proposition 2.4.11] on the subproblem in PPA which is (use also [11, Corollary 1, page 39])
@@
We compare the Algorithm 1 and the semismooth Newton method in [51] noted as PPA and SSN, respectively. Since the semismooth Newton method is also used for solving the subproblem in the PPA, we will not compare the iteration and wall-clock time in this paper. Only the optimal function values are compared in the following figures and tables.
@@
    We consider $N=2,5,10,20$ and for each value of NN we test 50 instances generated with different random parameters as described in Examples 27 and 28.
@@
In Table 2, we choose a less restrict comparison standard for Example 28 since the SSN hardly converges for most of the samples. NaN indicates that the median cannot be computed since most of the runs did not converge.
@@
In several examples, we observe that SSN converges with oscillatory behavior while PPA does not (see, for instance, Figures 2 and 3 which correspond to one instance selected from our experiments of Examples 27 and 28, respectively).
@@ 6 Conclusions
We contribute to the extension of the remarkable properties of the PPA from (strongly) convex functions to a broader class of functions, known as (strongly) quasar-convex. By generalizing the standard properties of proximity operators, we prove that PPA generates a minimizing sequence for quasar-convex functions and converges linearly in the strongly quasar-convex case, achieving a rate matching that of the strongly convex case. Additionally, we present several noteworthy examples of nonsmooth (strongly) quasar-convex functions, which have practical applications in machine learning, microeconomic theory, and beyond.
@@
Acknowledgements This research was partially supported by ANID–Chile under project Fondecyt Regular 1241040 (Lara) and Conselho Nacional de Desenvolvimento Científico e Tecnológico CNPq 153172/2024-0 (Liu).

### v2

@@ 5 Numerical Experiments
In this section, we illustrate that, as in the convex case, solving a sequence of proximal subproblems may improve the conditioning of the problem and thus its numerical behavior.
In the examples below, we consider classes of functions which are neither convex nor quasiconvex, since their sublevel sets are not convex and therefore represent hard problems for first-order algorithms.
@@ 5.1 Implementations
To solve the above two examples in this section via PPA, we cannot use the common smooth first-order method or second-order method for the subsproblems since both subproblems involved in Examples 28 and 29 are nonsmooth. Therefore, we use the Clarke necessary optimality condition [11, Proposition 2.4.11] on the subproblem in PPA which is (use also [11, Corollary 1, page 39])
@@
We compare Algorithm 1 with the semismooth Newton method in [51], referred to as PPA and SSN, respectively. Our comparison primarily focuses on the optimal function values attained by the two methods. This choice is motivated by the fact that the semismooth Newton method is employed as a subroutine within PPA, so that both methods share a common computational core, making direct comparisons in terms of iteration counts and computational time less informative.
Nevertheless, for completeness, we also report the wall-clock time. The numerical results show that PPA exhibits significantly better robustness and convergence behavior than SSN. On the other hand, when both methods successfully converge, PPA typically requires more computational time, reflecting the additional cost of solving subproblems at each iteration.
@@
    We consider $N=2,5,10,20$ and for each value of NN we test 50 instances generated with different random parameters as described in Examples 28 and 29.
@@
In Table 2, we choose a less restrictive comparison standard for Example 29 since the SSN hardly converges for most of the samples. NaN indicates that the median cannot be computed since most of the runs did not converge.
Table 3: Wall-clock time (s) of PPA and SSN for Example 28 (successful cases only).
Method $N=2$ $N=5$ $N=10$ $N=20$
SSN 0.4784 0.5375 0.5319 0.5288
PPA 15.94 18.46 15.99 12.28
Table 4: Empirical linear contraction ratios $R_{i}$ and $R_{f}$ for PPA, reported at iterations k = 1, 2, 3, 4, 5, 6, 7 and 8. The values remain uniformly below 1, corroborating linear convergence of PPA for Example 28.
Iteration 11 22 33 44 55 66 77 88
$R_{i}$ 0.1535 0.1721 0.2758 0.3713 0.3720 0.3720 0.3720 0.3720
$R_{f}$ 0.0332 0.0272 0.0290 0.1487 0.1384 0.1384 0.1384 0.1384
@@
In several examples, we observe that SSN converges with oscillatory behavior while PPA does not (see, for instance, Figures 2 and 3 which correspond to one instance selected from our experiments of Examples 28 and 29, respectively).
@@
  3. 3.
    For completeness, we report the wall-clock time of PPA and SSN for selected instances in Example 28. This comparison is restricted to these instances, since SSN fails to converge for many cases in Example 29, as indicated in Table 2. When both methods successfully converge, Table 3 shows that PPA is considerably more time-consuming than SSN. This is of course expected, as each iteration of PPA requires solving a subproblem via the semismooth Newton method, leading to a higher overall computational cost despite its improved robustness.
  4. 4.
    Table 4 reports the empirical ratios
    $R_{f}(k)=:\frac{f_{k+1}}{f_{k}},~~~~R_{i}(k)=\frac{\|x^{k+1}-x^{*}\|}{\|x^{k}-x^{*}\|},$
    generated by PPA for Example 28. All ratios stabilize strictly below one, providing clear numerical evidence of linear convergence of the function values and iterates, in agreement with the theoretical predictions for strongly quasar-convex function.
@@ 6 Conclusions
We contributed to the extension of the remarkable properties of PPA from (strongly) convex functions to a broader class of functions, known as (strongly) quasar-convex. By generalizing the standard properties of proximity operators, we prove that PPA generates a minimizing sequence for quasar-convex functions and converges linearly in the strongly quasar-convex case, achieving a rate matching that of the strongly convex case. Additionally, we present several noteworthy examples of nonsmooth (strongly) quasar-convex functions, which have practical applications in machine learning, microeconomic theory, and beyond.
@@
7 Declarations
7.1 Availability of Supporting Data
No datasets were generated during the present study. The Matlab code for the numerical experiments is available upon reasonable request from all authors.
7.2 Competing Interests
There are no conflicts of interest or competing interests related to this manuscript.
7.3 Funding
This research was partially supported by ANID–Chile under project Fondecyt Regular 1241040 (Lara) and Conselho Nacional de Desenvolvimento Científico e Tecnológico CNPq 153172/2024-0 (Liu).